OpenStudy (firejay5):

Simplify Each Expression. Show work and explain it. 32. m over m^2 - 4 + 2 over 3m + 6

OpenStudy (firejay5):

Well on #32, I am having trouble multiplying the denominators to make them the same

OpenStudy (anonymous):

"m over m^2 - 4 + 2 over 3m + 6" \[ \frac{m}{m^2-4}+\frac{2}{3m+6} \]

OpenStudy (firejay5):

why

OpenStudy (anonymous):

Just use this: \[ \frac{a}{b}+\frac{c}{d}=\frac{ad+cb}{bd} \]

OpenStudy (anonymous):

I think the + is separating two fractions.

OpenStudy (firejay5):

yes

OpenStudy (firejay5):

yea

OpenStudy (firejay5):

can you do it with me and then I'll do 33.) my self

OpenStudy (firejay5):

what do you mean?

OpenStudy (firejay5):

Equation editor

OpenStudy (firejay5):

can we get back to the problem

OpenStudy (firejay5):

or not @Hero

OpenStudy (firejay5):

the example he gave isn't going nowhere for me

OpenStudy (anonymous):

let \(b=m^2-4\) and \(d=3m+6\)

OpenStudy (anonymous):

\[ \frac{m}{m^2-4}+\frac{2}{3m+6}=\frac{(m)(3m+6)+(2)(m^2-4)}{(m^2-4)(3m+6)} \]

OpenStudy (anonymous):

@Firejay5 Do you see the pattern?

OpenStudy (firejay5):

Yes? Kind of how did you get that

OpenStudy (firejay5):

I need full out steps, like getting the denominators equal, etc.

OpenStudy (firejay5):

@Mertsj could you help with that

mathslover (mathslover):

hmn . \[\frac{m}{m^2-4}+\frac{2}{3m+6}\] Well I think that it would be Ok to factorize both denominators

OpenStudy (mertsj):

Yes. mathslover will be able to show you.

OpenStudy (firejay5):

@Mertsj you sure

mathslover (mathslover):

\(\color{blue}{m^2-4}\) is in the form of \(a^2-b^2\) \(a^2-b^2 = (a+b)(a-b)\) Use the above stated identity to factorize \(m^2-4 \) .

OpenStudy (mertsj):

yes

mathslover (mathslover):

Can you factor \(m^2-4\) @Firejay5 ?

OpenStudy (firejay5):

yes

mathslover (mathslover):

Can you tell me what would be the factorized form for \(x^2-4\) ?

OpenStudy (firejay5):

(x - 2) & (x + 2)

mathslover (mathslover):

Great! So \(x^2-4 = \color{blue}{(x+2)(x-2)}\) Now. we have the complete question as : \(\cfrac{m}{\color{red}{(m+2)(m-2)}}+\cfrac{2}{3m+6}\)

mathslover (mathslover):

Now take 3 common from 3m + 6. What you get?

OpenStudy (firejay5):

I did the other factorization as well

OpenStudy (firejay5):

3(m + 2)

mathslover (mathslover):

Great! \(\cfrac{m}{\color{red}{(m+2)(m-2)}}+\cfrac{2}{\color{orange}{3(m+2)}}\). Is what we have now.

mathslover (mathslover):

Now take LCM .

OpenStudy (firejay5):

m + 2

mathslover (mathslover):

Check it again

OpenStudy (firejay5):

and m - 2

mathslover (mathslover):

LCM of \(\cfrac{1}{a+b} + \cfrac{1}{2(a-b)} \) (denominators) is (a+b)(a-b)2

mathslover (mathslover):

Similarly do that there. we have \((x+2)(x-2) \quad\textbf{and}\quad 3(m+2)\) as the denominators

OpenStudy (firejay5):

it would be 3(m - 2) (m + 2)

mathslover (mathslover):

Yeah . and sorry there for using x on the place of m

mathslover (mathslover):

So I have : 3(m-2)(m+2) as LCM can you solve the fraction :

OpenStudy (firejay5):

so what would it be entered in equation editor

mathslover (mathslover):

I didn't get you,, please explain!

OpenStudy (firejay5):

type the problem in fraction form of what we have left

mathslover (mathslover):

hmn well it would be hard to say that I have : \(\cfrac{\textbf{yet to come}}{3(m+2)(m-2)}\)

mathslover (mathslover):

we have to solve numerator now

mathslover (mathslover):

can u do that?

OpenStudy (firejay5):

wouldn't the numerator m + 2

mathslover (mathslover):

No! See : \(\cfrac{1}{(a+b)} + \cfrac{2}{(a-b)} = \cfrac{1(a-b) + 2(a+b)}{(a-b)(a+b)}\)

OpenStudy (firejay5):

how does it work, you should + across

mathslover (mathslover):